# bibliography.bib

### v1

@@ 2 ReLU NN and free knot splines (FKS) as nonlinear function approximators
In this section, we define one-dimensional ReLU Neural Networks (NN) and free knot splines (FKS).
2.1 Deep and Shallow ReLU Neural Networks (NN)
We start by defining a standard feed forward neural network with a one-dimensional input $x_{0}\in\mathbb{R}$ and one-dimensional output $y_{NN}(x_{0},\bm{\theta})$ for some vector of parameters $\bm{\theta}$, described further below. We assume that this network has a width $W$ and a depth $L$ and uses a rectified linear unit (ReLU) $\sigma(x)=(x)_{+}$ as activation function which acts element wise on the input vector. Defining the network input by $x_{0}\in\mathbb{R}$, the first hidden layer is defined by
$\bm{x}_{1}=\sigma\left(\bm{A_{0}}x_{0}+\bm{b}_{0}\right)\in\mathbb{R}^{W}$
for $\bm{A}_{0},\bm{b}_{0}\in\mathbb{R}^{W}$ and the NN with $L$ hidden layers is then iteratively defined by
@@ 2.1 Deep and Shallow ReLU Neural Networks (NN)
$y(x,{\bm{\theta}})=\bm{A}_{L}\bm{x}_{L}+b_{L}\in\mathbb{R},$ (2)
with $\bm{A}_{L}\in\mathbb{R}^{1\times W}$ and $b_{L}\in\mathbb{R}$. In the simplest case we can consider a shallow ReLU network with width $W$ and depth $L=1$, otherwise the network is deep. The case $L=1$ is closest to standard approximation, but the real power of the NN comes from the expressivity seen when $L>1$ although as we shall see such networks are harder to train.
In this representation we define the set of parameters ${\bm{\theta}}=\{A_{k},\bm{b}_{k}\colon~{}k=0,\ldots,L\}$. For a given range of parameters ${\bm{\theta}}$ we define $\Gamma_{W,L}$ to be the set of all possible such functions $y(x)$. Observe that $\Gamma_{W,L}$ is not a linear vector space and that $A_{k}$ and $\bm{b}_{k}$ are not sparse in general.
We will apply this NN to approximate a function $u\colon[0,1]\to\mathbb{R}$ by finding an appropriate choice of $\bm{\theta}$ where we will not assume any regularity properties of $u$ at present. We assume that there exists a batch of training data $\{y_{x_{i}}\}_{i=0}^{s-1}\subset\mathbb{R}^{s}$ at training points $\{x_{i}\}_{i=0}^{s-1}\subset[0,1]$ of size $s$. Our aim is to use the NN so that the function $y(x)$ approximates the target $u(x)$ in the interval $[0,1]$. We use the standard means of training a NN to do this, including the use of a squared-error loss function and an optimiser such as Adam [KingBa15]. However, as we will see, crucial to the effectiveness of the training procedure is the use of a good initialisation to start the approximation, an effective loss function, and a careful preconditioning of the problem. We shall demonstrate this using the examples in this paper, and will also show how such a good initialisation, loss function and preconditioning can be constructed.
@@ 2.2 Linear spline approximations
Let $N\in\mathbb{N}$ be given and let
$10=k_{0}<k_{1}<\ldots k_{i}<\ldots<k_{N-1}=$ 1
be a set of $N$ knots in the interval $[0,1]$. For $i=0,\ldots,N-1$, we define a linear spline $\phi_{i}(x)$ to be the unique piecewise linear function so that
$\phi_{i}(k_{i})=1,\quad\mbox{and}\quad\phi_{i}(k_{j})=0\quad\mbox{if}\quad i\neq j.$
Each spline $\phi_{i}$ is weighted by some weight $w_{i}$, and we denote by
${\bm{\psi}_{FKS}}=\{(w_{i},k_{j})\colon i=0,\ldots,N-1,~{}j=1,\ldots,N-2\}$
the set of parameter values. Note that we exclude $k_{0}$ and $k_{N-1}$ in the definition of $\bm{\psi}_{FKS}$ as they are set as 0 and 1, respectively.
@@
In the usual case of piecewise linear approximation (used for example in the Finite Element method), the values of the knots $k_{i}$ are fixed and we refer to such approximations as either fixed knot piecewise linear approximations or regular approximations. The best function approximation is given by the conditioned problem of finding the optimal choice of weights $w_{i}$ for $i=0,\ldots,N-1$ and we denote by $\bm{\psi}_{PWL}=\{w_{i}\colon i=0,\ldots,N-1\}$ the associated parameter values.
@@
Note that (4) is a linear function of the coefficients $w_{i}$ and the set of all such possible functions is a linear vector space. The values of $w_{i}$ can be found simply by minimising the least squares loss. Equivalently, a linear matrix problem of the form $\bm{P}\bm{w}=\bm{q}$ for the coefficients $\bm{w}=(w_{i})_{i=0}^{N-1}\in\mathbb{R}^{N}$ can be solved, where the (well conditioned) sparse matrix $\bm{P}\in\mathbb{R}^{n\times n}$ and vector $\bm{q}\in\mathbb{R}^{n}$ are given. Whilst this is a simple process and widely used, for example in the Finite Element Method, such approximations lack expressivity, especially if the target function $u(x)$ is singular, or has natural length scales which are smaller than the smallest separation between the knots. The additional degrees of freedom given by choosing the knots $k_{i}$ in the FKS will (as we shall see) lead to a very significant increase in the accuracy of the approximation.
A distinguished subset of the set of free knot splines are the interpolating free knot splines where the set of knots $k_{i}$ for $i=1,\ldots,N-2$ can be freely chosen, but the weights are given by $w_{i}=u(k_{i})$ We denote the associated parameter values by $\bm{\psi}_{IFKS}=\{k_{i}\colon i=1,\ldots,N-2\}$.
@@
The interpolating free knot splines (FKS) $y(x,\bm{\psi}_{IFKS})=\Pi_{1}u(x)$ for the target function $u(x)$ is given by
$y(x,\bm{\psi}_{IFKS})\equiv\Pi_{1}u(x)=\sum_{i=0}^{N-1}u(k_{i})\phi_{i}(x),$ (5)
@@
The interpolating FKS is easier to construct than the best FKS approximation for $u(x)$ as the coefficients $w_{i}$ of the FKS are given directly by $w_{i}=u(k_{i}).$ It is thus not as expressive as the best FKS, but has (as we will see) the same rate of convergence as $N\to\infty$ for classes of singular function, and it is a much better approximation in general than a fixed knot piecewise linear approximation. Observe that finding the interpolating FKS still requires solving a nonlinear problem for the knots. We will make extensive use of this function as a way of both assessing the accuracy of a more general approximation to $u$ and of guiding the approximation procedure. We shall demonstrate that $\Pi_{1}y$ is a very good function to take as the initial start of the optimisation procedure for finding either the best ReLU NN approximation or the best FKS approximation to $u$. Indeed, if the optimal set $K$ of knot points is constructed for $\Pi_{1}u$ then this appears from calculations to be very close to the optimal set of knot points for the FKS, and equivalently the breakpoints of the ReLU NN representations.
3 Relations between the ReLU NN and the FKS representations and the associated conditioniing of the training problem
Any ReLU NN gives a function $y(x)$ which is piecewise linear, and is smooth (linear) everywhere apart from at a set of breakpoints where it changes gradient and has undefined curvature. These points are equivalent to the knot points $k_{i}$ of the FKS if they lie in the range of values of $x\in[0,1]$ over which the target function $u(x)$ is defined. In this sense the ReLU NN and the FKS are formally equivalent. However the map between ${\bm{\theta}}$ and $\{k_{i}\colon i=0,\ldots,N-1\}$ is complex and is related to the solution of a system of algebraic equations [DeVore_Hanin_Petrova_2021]. More precisely, the authors show that a NN of width $W$ and depth $L$ is formally equivalent to a FKS of size $N$ (where $N$ is the number of breakpoints) for $N<(W+1)^{L}$, and hence they can have the same expressivity, even though this is rarely seen in practice. Due to the linear dependencies between the linear pieces, the NN space $\Gamma_{W,L}$ is much smaller than the FKS space $\Sigma_{N,1}$ with $N$ satisfying $N<(W+1)^{L}$.
@@ 3.1 Equivalence of shallow ReLU NN and the FKS representations
Both a ReLU NN and an FKS are piecewise linear functions of $x$ and hence we can use the methodology of training one to inform the training of the other. To do this we now show how the coefficients of a shallow ReLU and FKS are related. For a shallow network with $L=1$ and $N=W$, we have breakpoints in the ReLU representation, which are equivalent to knots in the FKS at
@@
where $(\bm{b}^{0})_{i}$ and $(\bm{A}_{0})_{i}$ denote the $i$th components of vectors $\bm{b}_{0}$ and $\bm{A}_{0}$. For any ordered set of knots $k_{i}$, with $10=k_{0}<k_{1}<\ldots<k_{N-2}<k_{N-1}=$ 1 we can represent the piece-wise linear basis functions $\phi_{i}(x)$ of the FKS using the formula presented in [DeVore_Hanin_Petrova_2021] so that
$\phi_{i}(x)=\alpha_{i}ReLU(x-k_{i-1})-\beta_{i}ReLU(x-k_{i})+\gamma_{i}ReLU(x-k_{i+1})$ (7)
@@
Similarly
$\phi_{0}(x)=ReLU(k_{1}-x)/k_{1},\quad\phi_{N-1}(x)=ReLU(x-k_{N-2})/(1-k_{N-2}).$
This implies that FKS representation in terms of the basis functions can also be giveen in terms of ReLU functions as follows
$y(x,\bm{\psi}_{FKS})=\sum_{i=0}^{N-1}w_{i}\phi_{i}(x)\equiv w_{0}ReLU(k_{1}-x)/k_{1}+\sum_{i=0}^{N-2}c_{i}ReLU(x-k_{i})$
where the coefficients $c_{i}$ are rational functions of the knots $k_{i}$ and weights $w_{i}$. Taking careful note of the representation of the basis functions at the two ends of the interval we obtain:
$\begin{split}c_{0}&=\alpha_{1}w_{1},\\ c_{1}&=\gamma_{1}w_{2}-\beta_{1}w_{1},\\ c_{i}&=\gamma_{i}w_{i+1}-\beta_{i}w_{i}+\alpha_{i}w_{i-1},\quad i=2,\ldots,N-2.\end{split}$ (9)
Hence the two representations of the functions by the ReLU NN and by FKS are formally equivalent, with the equivalence of the ReLU breakpoints with knot points, and with the tri-diagonal linear map (9) between the weights ${\mathbf{w}}$ and scaling constants ${\mathbf{c}}$. Hence (in principle) both networks are equally expressive. However, the information from the respective parameter sets ${\bm{\theta}}$ and ${\bm{\psi}_{FKS}}$ is encoded differently. and is critical for their training and in particular for the well-posedness of the solution representation and the associated training algorithm. In general, the coefficients $w_{i}$ of the FKS are ${\cal O}(1)$, and may be monotone and of constant sign. However, if, for example, $k_{i}-k_{i-1}$ is small (which is often the case if we wish to approximate a function with a high gradient or high curvature) then the coefficients in (8) may well be large, and of non-constant sign, even if $u(x)$ has constant sign. As a consequence, the ReLU NN representation of a function may be ill-conditioned, even if the FKS representation is well-conditioned.
@@
For the target function $u(x)=x(1-x)$ for which the optimal set of approximating knots is uniform with spacing $h$, i.e. $h=k_{i+1}-k_{i}$, we have $c_{0}\approx 1$ and $c_{i}\approx-h$ for $i>0$. Observe that there is a large jump from $c_{0}$ to $c_{1}$ leading to a poor balancing of the ReLU NN representation, while the FKS representation is better balanced. If, instead, we assume that the knot points have a density of $1/|u^{\prime\prime}|$ then the values of $c_{i}$ are approximately constant and the ReLU NN representation is evenly balanced. This observation links the coefficients of the ReLU NN to the equidistribution condition we investigate in Section 4.2.
3.2 Conditioning of the training of the shallow ReLU NN and FKS representations
Using the results of the previous sub-section we now compare the conditioning of the problem of calculating the scaling factors $c_{i},i=0\ldots N-2$ and the weights $w_{i},i=1\ldots N-1$ of the shallow ReLU NN and the FKS representations respectively (to give the best $L_{2}$ approximation of the function $u(x)$) on the assumption that the knot locations $k_{i}$ are known. We show that the FKS problem (with the localised support basis functions) is well conditioned for all $N$ whereas the ReLU NN problem (for which basis functions have global support) is increasingly ill-conditioned for larger values of $N$.
It follows from (9) that
$\bm{c}=L\bm{w}$ (10)
for the tri-diagonal linear operator $L$ and we can compare the conditioning of the two problems by studying the properties of the matrix $L$. This matrix has a specific structure which we can exploit. It follows from (9) that
$c_{i}=\frac{w_{i+1}-w_{i}}{k_{i+1}-k_{i}}-\frac{w_{i}-w_{i-1}}{k_{i}-k_{i-1}}$ (11)
for $i=2,\ldots,N-2.$ Hence, the vectors $\bm{c}=(c_{i})_{i=1}^{N-2},{\mathbf{w}^{*}}=(w_{i})_{i=1}^{N-2}\in\mathbb{R}^{N-2}$ are related via
${\mathbf{c}}=T{\mathbf{w}}^{*}+\gamma_{N-2}w_{N-1}{\mathbf{e}}_{N-2},$ (12)
where ${\mathbf{e}}_{N-2}=(0,0,0,\ldots 1)^{T}\in R^{N-2}$ and $T\in R^{(N-2)\times(N-2)}$ is the symmetric tri-diagonal matrix with ${\mathbf{\beta}}_{i}$ on the leading diagonal, and ${\mathbf{\alpha}_{i}}$ and ${\mathbf{\gamma}_{i}}$ on the upper and lower first diagonals. Observe that to find ${\mathbf{w}}$ from ${\mathbf{c}}$ (and hence to invert the linear operator $L$) we can let $w_{N-1}$ be initially unknown, invert $T$ using the Thomas algorithm, to find ${\mathbf{w}}^{*}$ in terms of $w_{N-1}$ and then fixing $w_{N-1}$ from the relation
$w_{1}=c_{0}/\alpha_{1}.$
Evidently the conditioning of $L$ is then completely determined by the conditioning of $T$.
3.3 Uniformly spaced knots
We can make very precise estimates of the conditioning of the problem of finding ${\mathbf{w}}$ and ${\mathbf{c}}$ in this case. Suppose that we have uniformly spaced knots $k_{i}=i/(N-1)$, then (11) reduces to
$c_{i}=(N-1)(w_{i+1}-2w_{i}+w_{i-1})$
The linear operator $T$ then becomes the tri-diagonal Toeplitz matrix
$T=(N-1)\begin{pmatrix}-2&1&0&\cdots&\cdots&\cdots&\cdots&0\\ 1&-2&1&0&&&&\vdots\\ 0&1&-2&1&\ddots&&&\vdots\\ \vdots&0&\ddots&\ddots&\ddots&\ddots&&\vdots\\ \vdots&&\ddots&\ddots&\ddots&\ddots&0&\vdots\\ \vdots&&&\ddots&1&-2&1&0\\ \vdots&&&&0&1&-2&1\\ 0&\cdots&\cdots&\cdots&\cdots&0&1&-2\\ \end{pmatrix}\in\mathbb{R}^{N-2\times N-2}.$
$\lambda_{k}=-2(M-1)+2(M-1)\cos(\pi k/(M+1))\in(-4(N-1),0)$
for $k=1,\ldots,M$. In particular, we have $|\lambda_{1}|\leq\ldots\leq|\lambda_{M}|$, and for large $M$ we then have
$\lambda_{1}$ $=-2(M-1)+2(M-1)\cos\left(\frac{\pi}{M+1}\right)$
Similarly, for large $M$,
$\lambda_{M}$ $=-2(M-1)+2(M-1)\cos\left(\frac{M\pi}{M+1}\right)$
$\approx-2(M-1)+2(M-1)\left(-1+\frac{\pi^{2}}{2(M+1)^{2}}\right)=-4(M-1)-\frac{\pi^{2}(M-1)}{(M+1)^{2}}\approx-4(M-1),$
implying that the condition number $\kappa(T)$ satisfies
For the FKS representation of the approximation to the target function, we aim to solve the problem for the weights ${\mathbf{w}}$ given by
$M\bm{w}=\bm{u},$
where $M_{i,j}=\langle\phi_{i},\phi_{j}\rangle$, giving
$M=\frac{1}{6(N-1)}\begin{pmatrix}4&1&0&\cdots&\cdots&\cdots&\cdots&0\\ 1&4&1&0&&&&\vdots\\ 0&1&4&1&\ddots&&&\vdots\\ \vdots&0&\ddots&\ddots&\ddots&\ddots&&\vdots\\ \vdots&&\ddots&\ddots&\ddots&\ddots&0&\vdots\\ \vdots&&&\ddots&1&4&1&0\\ \vdots&&&&0&1&4&1\\ 0&\cdots&\cdots&\cdots&\cdots&0&1&4\\ \end{pmatrix}\in\mathbb{R}^{N\times N}.$
$6(N-1)\mu_{k}=4+2\cos(\pi k/(N+1))\in(2,6)\quad k=1,\ldots,N.$
Note that $\mu_{N}\leq\ldots\leq\mu_{1}$ with $26(N-1)\mu_{N}>$ 2, $66(N-1)\mu_{1}<$ 6 and condition number $\kappa(M)=\mu_{1}/\mu_{N}<3$. As $N\to\infty$ we have $6(N-1)\mu_{N}\to 2$, $6(N-1)\mu_{1}\to 6$ and
$\kappa(M)\to 3.$
The associated normal equations for ${\mathbf{w}}$ are given by
$M^{T}M\bm{w}=M^{T}\bm{u}.$
The boundedness of $\kappa(M)$ in the limit of large $N$ lies at the heart of the regularity, and eases the construction of the FKS approximation from the normal equations in this limit.
We now consider the related problem of finding the coefficients ${\mathbf{c}}$ of the ReLU approximation. This is of course closely linked to the problem of finding the FKS coefficients ${\mathbf{w}}$. For simplicity we will consider the case where $w_{0}=w_{N-1}=0$ so that the operators $L$ and $T$ are the same. From (10), we then obtain
$M^{T}MT^{-1}\bm{c}=M^{T}\bm{u}$
and multiplication by $(T^{-1})^{T}$ yields
$\tilde{M}^{T}\tilde{M}\bm{c}=\tilde{M}^{T}\bm{u}$
for the matrix
$\tilde{M}=MT^{-1}.$
To calculate the condition number of the matrix $\tilde{M}$ we note, that as the matrices $T$ and $M$ have exactly the same tri-diagonal Toeplitz structure, it follows from standard theory that they have identical eigenvectors ${\mathbf{e}}_{k}$, $k=1\ldots N$ with the eigenvector ${\mathbf{e}}_{k}$ having corresponding eigenvalues $\lambda_{k}$ and $\mu_{k}$ for the matrices $T$ and $M$ respectively. It follows immediately, that the eigenvectors of the matrix $\tilde{M}$ are also given by the vectors ${\mathbf{e}}_{k}$ with corresponding eigenvalues
$\nu_{k}=\mu_{k}/\lambda_{k}.$
Observe that for large $N$ we have $\nu_{1}\to-6N/\pi^{2}$ and $\nu_{N}\to-1/2N.$ Hence we deduce that as $N\to\infty$
$\kappa(\tilde{M})=\nu_{1}/\nu_{N}\to\frac{12\;N^{2}}{\pi^{2}}.$
In particular, this shows that $\kappa(\tilde{M})=\mathcal{O}(N^{2})$ for $N\gg 1$. When solving the normal equations for the weights ${\bm{c}}$ of the ReLU network, we must consider the condition number $\kappa$ of the matrix $\tilde{M}\tilde{M}^{T}$ which from the above calculation satisfies $\kappa={\cal O}(N^{4})$. This is very large even for moderate values of $N$ such as $N=64$, and is huge compared with the ${\cal O}(1)$ condition number for the FKS approximation. Classical optimisation methods (such as BFGS) have convergence rate dependent on
$(\sqrt{\kappa}-1)/(\sqrt{\kappa}+1)=1-{\cal O}(1/N^{2})$
This leads to very slow convergence for large $N$. We see presently that the performance of the Adam optimiser is similarly very slow on this problem.
3.4 Non-uniformly spaced knots
Hence (consistent with the estimate on a uniform mesh)
$\kappa(M)<3\;\frac{\max(k_{i+i}-k_{i-1})}{\min(k_{i+i}-k_{i-1}).}$
Note that this estimate for the condition number is smallest for uniform knots, and increases if the knot spacing is non-uniform.
In contrast, each row of the matrix $T$ has the respective coefficients:
Observe that $\alpha_{i}-\beta_{i}+\gamma_{i}=0$ for each $i$. It follows that the matrix $T$ is near singular; indeed if ${\mathbf{f}}=(1,1,1,1,,\ldots 1)^{T}$ then
$T{\mathbf{f}}={\mathbf{g}}=(\gamma_{1}-\beta_{1},0,\ldots,0,\alpha_{N-2}-\beta_{N-2})^{T}.$
Hence
$\|T^{-1}\|>\|f\|/\|g\|\to\infty,\quad\mbox{as}\quad N\to\infty.$
The condition number of $T$, and hence of $MT^{-1}$ thus increases without bound as $N$ increases. Numerical experiments strongly indicate that, as in the case of the matrix $M$, the condition number increases more rapidly for non-uniform knots than uniform knots.
4 Loss functions and the associated training methods

### v2

@@ 2 ReLU  NN and free knot splines (FKS) as nonlinear function approximators
In this section, we define univariate ReLU Neural Networks (NN) and free knot splines (FKS).
2.1 ReLU Neural Networks (ReLU NN)
We start by defining a standard feed forward neural network with a one-dimensional input $x_{0}\in\mathbb{R}$ and one-dimensional output $y(x_{0},\mathbf{\theta})$ for some vector of parameters $\mathbf{\theta}$, described further below. We assume that this network has a width $W$ and a depth $L$ and uses a rectified linear unit (ReLU) $\sigma(x)=(x)_{+}$ as activation function which acts element wise on the input vector. Defining the network input by $x_{0}\in\mathbb{R}$, the first hidden layer is defined by $\bm{x}_{1}=\sigma\left(\bm{A_{0}}x_{0}+\bm{b}_{0}\right)\in\mathbb{R}^{W}$ for $\bm{A}_{0},\bm{b}_{0}\in\mathbb{R}^{W}$ and the NN with $L$ hidden layers is then iteratively defined by
@@ 2.1 ReLU Neural Networks (ReLU NN)
$y(x,{\mathbf{\theta}})=\bm{A}_{L}\bm{x}_{L}+b_{L}\in\mathbb{R},$ (4)
with $\bm{A}_{L}\in\mathbb{R}^{1\times W}$ and $b_{L}\in\mathbb{R}$. In the simplest case we can consider a shallow ReLU  network with width $W$ and depth $L=1$, otherwise the network is deep. The case $L=1$ is closest to standard approximation, but the power and expressivity of the NN often arises when $L>1$, although such networks are harder to train and to analyse. In this representation we define the set of parameters ${\mathbf{\theta}}=\{A_{k},\bm{b}_{k}\colon~{}k=0,\ldots,L\}$. For a given range of parameters ${\mathbf{\theta}}$ we define $\Gamma_{W,L}$ to be the set of all functions $y(x)$ which can be represented by the NN. Observe that $\Gamma_{W,L}$ is not a linear vector space and that $A_{k}$ and $\bm{b}_{k}$ are not sparse in general. We apply this NN to approximate a function $u(x)\colon[0,1]\to\mathbb{R}$ by finding an appropriate choice of $\mathbf{\theta}$. We assume that there exists a batch of training data $\{y_{x_{i}}\}_{i=0}^{s-1}\subset\mathbb{R}^{s}$ at training points $\{x_{i}\}_{i=0}^{s-1}\subset[0,1]$ of size $s$. Typically this uses the standard means of training a NN to do this, including the use of a squared-error $L_{2}$ loss function given by $L=\|y-u\|_{2}^{2}$ and an optimiser such as Adam [17]. However, as we will see, crucial to the effectiveness of the training procedure is the use of a good initialisation to start the approximation, an effective (convex) loss function, and a careful preconditioning of the problem. We shall demonstrate this using the examples in this paper, and will also show how an effective two-level training method can be constructed, using a suitably convex loss function combined with preconditioning.
@@ 2.2 Linear spline approximations
Let $N\in\mathbb{N}$ be given and let $10=k_{0}<k_{1}<\ldots k_{i}<\ldots<k_{N-1}=$ 1 be a set of $N$ knots in the interval $[0,1]$. For $i=0,\ldots,N-1$, we define a linear spline $\phi_{i}(x)$ to be the unique piecewise linear function so that $\phi_{i}(k_{i})=1$, and $\phi_{i}(k_{j})=0$ if $i\neq j.$ Each spline $\phi_{i}$ is weighted by some weight $w_{i}$, and we denote by ${\mathbf{\psi}_{\rm FKS}}=\{(w_{i},k_{j})\colon i=0,\ldots,N-1,~{}j=1,\ldots,N-2\}$ the set of parameter values. Note that we exclude $k_{0}$ and $k_{N-1}$ in the definition of $\mathbf{\psi}_{\rm FKS}$ as they are set as 0 and 1, respectively.
@@
In the more usual case of piecewise linear approximation, the values of the knots $k_{i}$ are fixed and we refer to such approximations as either fixed knot piecewise linear approximations or regular approximations. The best function approximation is given by the well-conditioned problem of finding the optimal choice of weights $w_{i}$ for $i=0,\ldots,N-1$ and we denote by $\mathbf{\psi}_{\rm PWL}=\{w_{i}\colon i=0,\ldots,N-1\}$ the associated parameter values.
@@
Note that (6) is a linear function of the coefficients $w_{i}$ and the set of all such possible functions is a linear vector space. The values of $w_{i}$ can be found simply by minimising the least squares loss. Equivalently, a linear matrix problem of the form $\bm{P}\bm{w}=\bm{q}$ for the coefficients $\bm{w}=(w_{i})_{i=0}^{N-1}\in\mathbb{R}^{N}$ can be solved, where the (well conditioned) sparse matrix $\bm{P}\in\mathbb{R}^{n\times n}$ and vector $\bm{q}\in\mathbb{R}^{n}$ are given. Whilst this is a simple process and widely used, for example in the Finite Element Method, such approximations lack expressivity, especially if the target function $u(x)$ is singular, or has natural length scales which are smaller than the smallest separation between the knots. The additional degrees of freedom given by choosing the knots $k_{i}$ in the FKS leads to a very significant increase in the accuracy of the approximation.
A distinguished subset of the set of free knot splines are the interpolating free knot splines where the set of knots $k_{i}$ for $i=1,\ldots,N-2$ can be freely chosen, but the weights are given by $w_{i}=u(k_{i})$ We denote the associated parameter values by $\mathbf{\psi}_{\rm IFKS}=\{k_{i}\colon i=1,\ldots,N-2\}$.
@@
The interpolating free knot splines (FKS) $y(x,\mathbf{\psi}_{\rm IFKS})=\Pi_{1}u(x)$ for the target function $u(x)$ is given by
$y(x,\mathbf{\psi}_{\rm IFKS})\equiv\Pi_{1}u(x)=\sum_{i=0}^{N-1}u(k_{i})\phi_{i}(x),$ (7)
@@
The interpolating FKS is easier to construct than the best FKS approximation for $u(x)$ as the coefficients $w_{i}$ of the FKS are given directly by $w_{i}=u(k_{i})$ and only the knots need to be determined. It is thus not as expressive as the best FKS, but has the same rate of convergence as $N\to\infty$ for classes of singular functions [15], and it is a much better approximation in general (especially for smaller values of $N$) than a fixed knot piecewise linear approximation. Observe that finding the interpolating FKS still requires solving a nonlinear problem for the knots. We will make extensive use of this function as a way of guiding the approximation procedure and also assessing the accuracy of a more general FKS/ReLU NN approximation to $u$. In particular $\Pi_{1}y$ is a very good function to take as the initial start of the optimisation procedure for finding either the best ReLU NN approximation or the best FKS approximation to $u$. This is because the optimal set $K$ of knot points is constructed for $\Pi_{1}u$ appears from extensive calculations to be very close to the optimal set of knot points for the FKS, and equivalently the breakpoints of the ReLU NN representations.
3 Relations between the ReLU NN and the FKS representations
Any ReLU NN gives a function $y(x)$ which is piecewise linear, and is smooth (linear) everywhere apart from at a set of breakpoints where it changes gradient and has undefined curvature. These points are equivalent to the knot points $k_{i}$ of the FKS if they lie in the range of values of $x\in[0,1]$ over which the target function $u(x)$ is defined. In this sense the ReLU NN and the FKS are formally equivalent. For a deep NN the map between ${\mathbf{\theta}}$ and $\{k_{i}\colon i=0,\ldots,N-1\}$ is complex and is related to the solution of a system of algebraic equations [5]. More precisely, the authors show that a NN of width $W$ and depth $L$ is formally equivalent to a FKS of size $N$ (where $N$ is the number of breakpoints) for $N<(W+1)^{L}$, and hence they can have the same expressivity, even though this is rarely seen in practice. Due to the linear dependencies between the linear pieces, the NN space $\Gamma_{W,L}$ is usually much smaller than the FKS space $\Sigma_{N,1}$ with $N\ll(W+1)^{L}$. Accordingly we focus on the case of shallow NN where the correspondence is direct.
@@ 3.1 Equivalence of shallow ReLU  NN and the FKS representations
Both a ReLU  NN and an FKS are piecewise linear functions of $x$ and we can use the methodology of training one to inform the training of the other. For a shallow network with $L=1$ and $N=W$, we have breakpoints in the ReLUrepresentation, which are equivalent to knots in the FKS at
@@
where $(\bm{b}^{0})_{i}$ and $(\bm{A}_{0})_{i}$ denote the $i$th components of vectors $\bm{b}_{0}$ and $\bm{A}_{0}$. For any ordered set of knots $k_{i}$, with $10=k_{0}<k_{1}<\ldots<k_{N-2}<k_{N-1}=$ 1 we can represent the piece-wise linear basis functions $\phi_{i}(x)$ of the FKS using the formula presented in [5] so that
$\phi_{i}(x)=\alpha_{i}{\rm ReLU}(x-k_{i-1})-\beta_{i}{\rm ReLU}(x-k_{i})+\gamma_{i}{\rm ReLU}(x-k_{i+1})$ (9)
@@
Similarly, $\phi_{0}(x)={\rm ReLU}(k_{1}-x)/k_{1},\quad\phi_{N-1}(x)={\rm ReLU}(x-k_{N-2})/(1-k_{N-2}).$ This implies that FKS representation in terms of the basis functions can also be given in terms of ReLU functions as follows
$y(x,\mathbf{\psi}_{\rm FKS})=\sum_{i=0}^{N-1}w_{i}\phi_{i}(x)\equiv w_{0}{\rm ReLU}(k_{1}-x)/k_{1}+\sum_{i=0}^{N-2}c_{i}{\rm ReLU}(x-k_{i})$ (11)
where the ReLUcoefficients $c_{i}$ are rational functions of the knots $k_{i}$ and weights $w_{i}$ of the FKS. Taking careful note of the representation of the basis functions at the two ends of the interval we obtain:
$\begin{split}c_{0}&=\alpha_{1}w_{1},\quad c_{1}=\gamma_{1}w_{2}-\beta_{1}w_{1},\\ c_{i}&=\gamma_{i}w_{i+1}-\beta_{i}w_{i}+\alpha_{i}w_{i-1},\quad i=2,\ldots,N-2.\end{split}$ (12)
Hence the two representations of the functions by the ReLU NN and by FKS are formally equivalent, with the equivalence of the ReLUbreakpoints with knot points, and with the tri-diagonal linear map (12) between the weights ${\mathbf{w}}$ and scaling constants ${\mathbf{c}}$. Hence (in principle) both networks are equally expressive. However, the information from the respective parameter sets ${\mathbf{\theta}}$ and ${\mathbf{\psi}_{\rm FKS}}$ is encoded differently, and is critical for their training and in particular for the well-posedness of the solution representation and the associated training algorithm. In general, the coefficients $w_{i}$ of the FKS are ${\mathcal{O}}(1)$, and may be monotone and of constant sign. However, if, for example, $k_{i}-k_{i-1}$ is small (which is often the case if we wish to approximate a function with a high gradient or high curvature) then the coefficients in (10) may well be large, and of non-constant sign, even if $u(x)$ has constant sign. As a consequence, the ReLU NN representation of a function may be ill-conditioned, even if the FKS representation is well-conditioned.
3.2 Weights for the IFKS and equivalent ReLU  NN
@@
For the target function $u(x)=x(1-x)$ for which the optimal set of approximating knots is uniform with spacing $h$, i.e. $h=k_{i+1}-k_{i}$, we have $c_{0}\approx 1$ and $c_{i}\approx-h$ for $i>0$. Observe that there is a large jump from $c_{0}$ to $c_{1}$ leading to a poor balancing of the ReLU NN representation, while the FKS representation is better balanced. If, instead, we assume that the knot points have a density of $1/|u^{\prime\prime}|$ then the values of $c_{i}$ are approximately constant and the ReLU NN representation is evenly balanced. This observation links the coefficients of the ReLU NN to the equidistribution condition and we investigate this further in Section 5.
4 Training the ReLU NN and the FKS